# Random Utility with Aggregation

### v2

@@ Abstract
We characterize when discrete-choice datasets that involve aggregation, such as category-level items or an outside option, are consistent with a random utility model (RUM). The underlying alternatives that an aggregated category represents may differ across individuals and remain unobserved by the analyst. We characterize the observable implications of RUMs with unknown composition of aggregated categories and show that they are surprisingly weak, implying only limited monotonicity of choice frequencies and standard RUM consistency on unaggregated menus. These restrictions are insufficient to justify the aggregated random utility model (ARUM) commonly assumed in empirical work. We identify two sufficient conditions that restore the implication of ARUM: non-overlapping preferences and menu-independent aggregation. Simulations show that violations of these conditions generate estimation bias, highlighting the practical importance of how data are aggregated.
@@ 1 Introduction
In empirical studies of discrete choice, it is common practice to aggregate alternatives. For instance, in car sales data, all variants of the Toyota RAV4 despite differing in features and prices, are often grouped under a single category, “RAV4.” A particularly important form of aggregation involves the outside option. In many empirical IO studies (e.g., Nevo (2001); Shum (2016)), the outside option is modeled as a single alternative that aggregates all unlisted choices. For example, in Nevo (2001), which studies consumer demand for 25 cereal brands, the market share of the outside good is defined as the residual: the difference between one and the total market share of the inside goods, implicitly capturing all other breakfast options, such as pancakes or omelettes. [^2] See Appendix A: Data, page 337.
The standard modeling approach assumes a random utility model (RUM) defined directly over the aggregated categories—what we call the aggregated random utility model (ARUM). That is, researchers assume that each category corresponds to a single alternative, and that preferences are defined over these categories as if they were atomic options.
However, this modeling simplification obscures an important complication: while the analyst only observes aggregate frequencies, the consumer may evaluate the underlying alternatives composing an aggregated category very differently. In the car example, different versions of the RAV4 such as the Gasoline LE and the Plug-in Hybrid XSE may vary significantly in both desirability and prices. Likewise, the outside option may encompass alternatives that differ greatly in price and quality. Furthermore, the exact composition of an aggregated category may be unobservable to the analyst. As a result, observed choice frequencies over aggregated categories may not represent consumers’ underlying preferences.
Our paper is the first to formally analyze the implications of unknown composition in aggregated categories. We characterize the observable restrictions of RUM in this setting and show precisely how they are weaker than those implied by ARUM. Through a simulation exercise, we further demonstrate that this weakness can lead to significant estimation biases. Building on these findings, we provide theoretical results that offer guidance on how to define aggregated categories in ways that mitigate such biases.
@@ 1.1 Preview of theoretical results
To illustrate our contribution, consider a similar setup that was studied in Nevo (2001), which involves 25 cereal brands along with an outside good that contains all other breakfast options. Importantly, the composition of the outside good may differ across individuals in the population. For instance, some consumers may not consider omelette a viable option due to egg shortages resulting from bird flu, while others may rule out pancakes because essential ingredients such as syrup are unavailable in stores.
We formalize this using two key concepts. An aggregation function $X$ specifies the possible underlying goods that an aggregated category may represent—for example, the outside option may include omelettes, pancakes, or other breakfast items. A composition distribution $\lambda$ then assigns, for each menu, probabilities to these possible realizations of the outside option. Thus, when the outside option is present, $\lambda$ determines whether a consumer faces only omelettes, only pancakes, both, or some other combination of breakfast goods. Importantly, analysts observe only the reduced choice frequencies over categories, not the aggregation function or the composition distribution.
We say that the observed dataset $\rho$ is consistent with a RUM if there exists (i) a composition distribution $\lambda$ with an aggregation function $X$ and (ii) a probability distribution $\mu_{{\scriptscriptstyle\mathscr{X}}}$ over rankings (i.e., linear orders) $\succ$ on the set $\mathscr{X}$ of the underlying alternatives (in this case, the 25 cereal brands along with other breakfast options) [^3] We write $\mu_{{\scriptscriptstyle\mathscr{X}}}$ to clarify that the measure is defined on rankings over the underline alternatives $\mathscr{X}$, not the aggregated categories., such that for all subsets $D$ of cereals and $x\in D$, the market share $\rho(D\cup\text{out},x)$ of cereal $x$ in a market where the set $D$ of cereals is available equals
@@
Our first question is: What are the observable implications for $\rho$ that are consistent with some pair $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$ as described above? We show that these implications are characterized exactly by two conditions. The first is Limited Monotonicity—that is, the choice frequencies on menus containing only non-aggregated categories decrease when aggregated categories are added to the choice set. The second is a condition that we call Partial RU-rationality, that is on the subset of choice sets that do not contain any aggregated categories, choices must be rationalizable by a standard RUM over categories. See Theorems 3.1 and 3.12 in Section 3 for the formal statements.
Crucially, these two conditions are insufficient to guarantee the existence of an aggregated random utility model (ARUM), that is, a distribution over preference rankings on aggregated categories. This is because ARUMs require much more structure. The simplest implication of ARUM is full monotonicity — that is, choice frequencies decrease when alternatives are added to any choice set. The RU-rational choice functions may violate full monotonicity when the choice set contains aggregated categories. This gap is especially concerning given how commonly ARUMs are used in empirical work.
In addition to the characterization in Theorem 3.1, Theorem 3.6 provides a complementary characterization of RU-rationality focusing on the behavior of each consumer in the population. It shows that RU-rationality allows behavior ruled out by ARUM: an agent may follow a well-defined ranking in some cases but default to the outside option in others. This can be interpreted as menu effects, where complex or unfamiliar menus lead the agent to choose the outside option. The result underscores that, without further restrictions, RU-rationality does not imply ARUM.
Given these results, we next ask: under what additional conditions can RU-rational stochastic choice functions also be rationalized by an ARUM? This question is practically important, as it guides how aggregated categories should be defined in empirical work.
We identify two sufficient conditions. The first is the non-overlapping condition on $\mu_{{\scriptscriptstyle\mathscr{X}}}$: all of the alternatives that underlie an aggregated category must occupy similar positions in consumers’ rankings. In the cereal example, this means no consumer places some cereals between other breakfast options such as omelettes and pancakes. When overlap occurs—e.g., some consumers prefer omelettes to certain cereals but not others—the condition fails (Proposition 4.2). The second is menu independence of the composition distribution $\lambda$: the distribution of underlying components must not vary with the choice set. In empirical IO, where each menu corresponds to a market, this requires that $\lambda$ remain constant across markets. For instance, if eggs are scarce in one market (making omelettes unavailable) but plentiful in another (so omelettes remain available), the condition is violated (Proposition 4.5).
@@ 1.2 Simulation results
Our results imply that when either condition—non-overlapping preferences or menu independence—fails, the observed choice frequencies diverge from those predicted by an ARUM. Estimating an ARUM (such as logit model over aggregated categories) can therefore generate substantial bias. Assessing this bias empirically is crucial for understanding the practical relevance of our theory.
We examine this prediction through simulations. We assume that the true choice dataset defined on the underlying alternatives $\mathscr{X}$ is a logit model with fixed utility levels. We then fix values of the composition distribution $\lambda$ and an aggregation function $X$ and construct the reduced dataset $\rho$ over aggregated categories. Finally, we re-estimate utilities from $\rho$ as if it followed a logit model over aggregated categories. When $\lambda$ is menu-independent or preferences are non-overlapping, $\rho$ is consistent with an ARUM and bias must be low, otherwise the bias should become large as $\rho$ cannot be represented by an ARUM.
As our measure of bias, we examined the difference in estimated utilities between two non-aggregated alternatives, and we also computed the geometric distance from $\rho$ to the set of ARUMs. Both measures increase as $\lambda$ becomes more menu-dependent or preferences more overlapping. In some cases, the bias was substantial enough to reverse the ranking of alternatives—for example, even when $u(x)>u(y)$ in the true model, the estimates produced $\hat{u}(y)>\hat{u}(x)$. These findings align with our theoretical predictions and underscore the empirical importance of potential biases. Further details are provided in Section 5.
@@ 1.3 Related Literature
There is no paper studying stochastic choice focusing on the unknown composition of aggregated categories. There are, however, several papers that are related in terms of mathematical setup or motivation.
@@
The first line of research related to our paper is the recent work on random attention models, such as Cattaneo, Ma, Masatlioglu, and Suleymanov (2020), Aguiar, Boccardi, Kashaev, and Kim (2023), Brady and Rehbeck (2016), Manzini and Mariotti (2014),and Horan (2019). See Strzalecki (2022) for a survey of the literature. These papers study models in which agents do not consider the full available choice set but instead randomly consider a subset. In particular, Horan (2019) studies the case where the choice frequency of the outside or “default” option is not observable under the random attention model.
The resemblance is superficial. The models of random attention assume that each decision maker randomly attends to a subset of the alternatives. Our setting, by contrast, is driven by aggregation-induced unobservability: the analyst sees aggregated categories whose internal composition differ across consumers and menus. This distinction has sharp mathematical consequences. In our setup, when an aggregated category is present, at least one of its underlying components is available to the consumer. The models of random attention impose no such requirement. This constraint yields testable implications for RUM in our paper.
@@
Azrieli and Rehbeck (2022) studies a mathematically related model but with a different focus. They characterize marginal stochastic choice—frequencies not conditioned on particular menus—assuming the analyst observes the distribution of menus but not within-menu choices. By contrast, in our framework the distribution over available alternatives is unobserved and captured by $\lambda$, while we observe reduced stochastic choice over aggregated categories. Although Azrieli and Rehbeck (2022) also allows for endogenous menu availability, their analysis centers on models such as Gul and Pesendorfer (2001) and Kreps (1979), and addresses rationalizability rather than estimation bias.
Apesteguia and Ballester (2016) also study aggregation of stochastic choice data, but with a different motivation. They examine whether aggregate choice behavior, obtained by combining individual choices, remains in the same class as the underlying individual behaviors. By contrast, our model aggregates alternatives and allows for aggregation across agents facing different choice problems.
@@
Our companion paper Kono, Saito, and Sandroni (2025) also studies the RUM with an outside option, but in a different setup. The paper does not analyze the case of multiple aggregated categories. Moreover, from the perspective of the observed dataset, the key distinction between the two papers is that the present study takes the reduced dataset $\rho$ as the observed object, whereas the companion paper treats the full dataset over the underlying alternatives (with missing information) as given. In addition, the companion paper assumes the outside option always corresponds to the full set of underlying alternatives—that is, $\lambda$ is menu-independent and degenerate. Even in this idealized case, much of the implications of the RUM are lost. In the present paper, we extend that framework by allowing $\lambda$ to be non-degenerate or menu-dependent, and we characterize the observable implications of the RUM under this more general specification.
@@
A few papers in the empirical IO literature such as Brownstone and Li (2017), Stafford (2018), and Huang and Rojas (2013) have examined potential biases arising from aggregation and misspecification of the outside option. More recently, Zhang (2023) analyzed biases resulting from misspecifying market size, and hence the market share of the outside good. However, all of these papers are purely empirical in nature and do not address the issue of unknown composition of aggregated categories or outside options theoretically.
@@
Although less directly related, classical statistical work has developed a comprehensive framework for making inferences from data that are partially observed or categorized. For example, see Heitjan (1989) for a review of grouped data; Blumenthal (1968) and Nordheim (1984) for categorized data; and Heitjan and Rubin (1990, 1991) for coarse data. In this framework, subsets of the sample space are observed rather than exact realizations. [^4] For example, if the sample space is the real line (e.g., blood pressure), we may only observe a random variable taking interval values (e.g., an interval of blood pressures containing the true value). A key distinction from our setting is that, in our model, uncertainty arises from the actual choice set an agent faces, rather than from the data collection or coarsening process.
@@ 2.1 Alternatives
Let ${\mathscr{A}}=\{a_{0},....,a_{N}\}$ be the (finite) set categories. Let $\mathscr{X}$ be the set of underlying alternatives. We assume that $\mathscr{X}$ is finite. [^5] To ensure the existence of all aggregation functions relevant to the RUM, we assume that $\mathscr{X}$ contains at least $|\mathscr{A}_{A}|^{2}$ distinct elements. We consider a partition $\{\mathscr{A}_{N},\mathscr{A}_{A}\}$ of $\mathscr{A}$. [^6] That is, $\mathscr{A}_{N}\cup\mathscr{A}_{A}=\mathscr{A}$ and $\mathscr{A}_{N}\cap\mathscr{A}_{A}=\emptyset$. We interpret $\mathscr{A}_{N}$ as the set of nonaggregated categories; $\mathscr{A}_{A}$ the set of aggregated categories.
@@
A function $X:\mathscr{A}\to 2^{\mathscr{X}}\setminus\{\emptyset\}$ is called an aggregation function if (i) $X(a_{i})\cap X(a_{j})=\emptyset$ for any distinct $i,j$; (ii) $X(a)$ is singleton if $a\in\mathscr{A}_{N}$; (iii) $X(a)$ is non-singleton if $a\in\mathscr{A}_{A}$.
We call $X$ an aggregation function because the function (or more precisely, its inverse) specifies how each underlying alternative is assigned to a category. Condition (i) requires that every underlying alternative be assigned to exactly one category. Conditions (ii) and (iii) justify the interpretation that $\mathscr{A}_{N}$ represents the set of non-aggregated categories, while $\mathscr{A}_{A}$ represents the set of aggregated categories.
In many parts of the paper, we consider the setup in which there is only one aggregated category $a_{0}$ called outside option. In the setup, we assume $\mathscr{A}_{N}=\{a_{0}\}$.
@@
(Cereal Choice) Consider a similar setup to Nevo (2001), who estimates demand over the 25 cereal brands. As mentioned in the introduction, a key feature of the analysis is the inclusion of an outside good, representing all breakfast options not explicitly listed among the 25 cereal brands in the dataset. In the setup, all available alternatives are

### v3

@@ Abstract
We study random utility (RU) rationality with aggregation when the underlying alternatives in each aggregate vary across consumers and are unobserved, as is typical for an outside option. RUM over the underlying alternatives is the natural assumption on the data generating process, while an aggregated random utility model (ARUM) is the standard empirical tool. We characterize RU rationality in three frameworks and show its testable implications are substantially weaker than those of an ARUM. We provide two independent conditions for their equivalence: non-overlapping preferences within aggregates and menu-independent aggregation. Simulations show that violating either condition produces meaningful estimation bias when imposing an ARUM.
@@ 1 Introduction
In empirical economics, it is common practice to aggregate alternatives. For instance, in meat sales data (e.g. Eales and Unnevehr (1982)), all beef products, despite differing in quality and price, are sometimes grouped under a single aggregate, “beef.” Another particularly important form of aggregation involves the outside option. A canonical treatment in the standard IO framework models the outside option as a single alternative that aggregates all otherwise unlisted choices (see, e.g., the textbook treatment in Shum (2016)). For instance, in Nevo (2001), which studies consumer demand for popular cereal brands, the market share of the outside good is defined as the residual: the difference between one and the total market share of the inside goods, implicitly capturing all other breakfast options, such as pancakes or omelettes. [^2] See Appendix A: Data, page 337 of Nevo (2001).
The standard modeling approach assumes a random utility model defined directly over the aggregates—what we call the aggregated random utility model (ARUM). That is, researchers assume that each aggregate corresponds to a single alternative, and that preferences are defined over these aggregates as if they were atomic.
However, this simplification obscures an important complication. While the analyst only observes choice frequencies over aggregates, consumers evaluate the underlying alternatives that compose each aggregate—alternatives that may differ substantially in desirability. In the meat example, ground beef and table cuts vary significantly in both features. Likewise, the outside option may encompass alternatives that differ greatly in quality and price. Thus, a random utility model (RUM) should in principle be defined over the underlying alternatives, not over analyst-defined aggregates. The difficulty is that the exact composition of an aggregate may be heterogeneous across consumers and unobservable to the analyst. For instance, some consumers may not consider omelettes a viable breakfast option due to egg shortages from bird flu, while others may rule out pancakes because essential ingredients are unavailable.
In this paper, we formalize a RUM that accommodates such heterogeneous and unknown composition of aggregates, and we study its relationship to the ARUM. This relationship is important because the two models play distinct roles. A RUM over the underlying alternatives is the natural assumption on the data generating process, since consumers evaluate and choose among the actual underlying goods available to them. An ARUM, by contrast, is an empirical tool that enables tractable estimation by treating each aggregate as a primitive object of choice. Whether this simplification is valid depends on whether the choice frequencies over aggregates, generated by a RUM on the underlying alternatives, are also consistent with an ARUM.
This paper is the first to characterize the testable implications of a RUM when aggregated categories have heterogeneous and unknown composition. Using three different frameworks, we show precisely how these implications are weaker than those of an ARUM. Through simulations, we demonstrate that this gap can lead to significant estimation biases when an ARUM is incorrectly imposed. Building on these findings, we provide theoretical guidance on how to define aggregated categories in ways that mitigate such biases.
@@ 1.1 Preview of theoretical results
To formalize our framework, consider a setup involving popular cereal brands along with an outside good that contains all other breakfast options. In this setup, the outside option is an aggregated alternative and all other alternatives are not aggregated. All definitions and results below are naturally generalized into the case of multiple aggregated alternatives.
The composition of the outside good may differ across individuals in the population due to local availability. We formalize this using two key concepts. An aggregation correspondence $X$ specifies the possible underlying goods that an aggregate may represent–for example, the outside option may include omelettes, pancakes, or other breakfast items. A composition distribution $\lambda$ then assigns, for each menu, probabilities to these possible realizations of the outside option. Thus, when the outside option is present, $\lambda$ determines whether a consumer faces only omelettes, only pancakes, both, or some other combination of breakfast goods. Crucially, the analyst does not observe the aggregation correspondence or the composition distribution.
We say that the choice frequencies $\rho$ is consistent with a RUM if there exists (i) a composition distribution $\lambda$ with an aggregation correspondence $X$ and (ii) a probability distribution $\mu_{{\scriptscriptstyle\mathscr{X}}}$ over rankings (i.e., linear orders) $\succ$ on the set $\mathscr{X}$ of the underlying alternatives (in this case, the popular cereal brands along with other breakfast options) [^3] We write $\mu_{{\scriptscriptstyle\mathscr{X}}}$ to clarify that the measure is defined on rankings over the underlying alternatives $\mathscr{X}$, not the aggregates., such that for all subsets $D$ of cereals and $x\in D$, the market share $\rho(D\cup\text{out},x)$ of cereal $x$ in a market where the set $D$ of cereals is available equals
@@
We theoretically study the implications of RU-rationality and compare them with those of ARUM along three dimensions. First, we characterize the implications of RU-rationality in terms of choice frequencies. Theorem 3.1 show that RU-rationality is equivalent to two conditions: (i) Limited Monotonicity—choice frequencies over non-aggregated alternatives weakly decrease when the aggregated alternatives are added to the menu; and (ii) Partial RU-rationality—on menus containing only non-aggregated alternatives, choices frequencies satisfy standard RUM restrictions. These conditions are substantially weaker than ARUM. Most notably, Limited Monotonicity does not require that choice frequencies decrease monotonically on menus already containing aggregated alternatives. This can occur because adding a new alternative may signal information about the composition of the aggregate alternatives—for instance, the presence of a premium cereal brand may indicate a high-income market where the outside option contains more attractive alternatives.
Second, we characterize the admissible deterministic behaviors under RU-rationality and compare them with those under ARUM. The deterministic behaviors completely characterize the RU-rational (respectively ARU-rational) stochastic choice functions in the sense that every stochastic choice function of either class can be written as a convex combination of deterministic behaviors. Under ARUM, admissible behaviors are precisely the rational choice functions induced by linear orders. Theorem 3.2 shows that RU-rationality admits a strictly richer class: an agent may maximize a linear order on some menus while defaulting to an aggregate such as the outside option on other menus. This can be interpreted as menu-dependent consideration—when faced with unfamiliar or complex choice sets, consumers may revert to the default rather than evaluate all options. Formally, the set of RU-rational stochastic choice functions forms a polytope whose vertices are these “menu-effect” choice functions, and the ARUM polytope is a strict subpolytope with double-exponentially fewer vertices.
Third, we characterize RU-rationality when the aggregation correspondence is known to the analyst but the composition distribution remains unknown. Theorem 3.3 establishes that as the number of underlying alternatives in each aggregated alternative grows, the set of RU-rationalizable choice functions expands and stabilizes once it exceeds a threshold. Beyond the threshold, RU-rationality is fully characterized by Limited Monotonicity and Partial RU-rationality. Theorem 3.4 shows that even below this threshold, the RU-rationalizable set may closely approximate the unconstrained case. These results suggest that in practice, Limited Monotonicity and Partial RU-rationality capture the empirically relevant implications of RUM; thus the implication of RU-rationality is significantly weaker than that of ARU-rationality.
Finally, given the gap between RU-rationality and ARU-rationality, we ask: under what conditions do they coincide? We identify two independent necessary and sufficient conditions. The first is non-overlapping preferences in the support of $\mu_{\mathscr{X}}$ (Proposition 4.1): if all underlying alternatives that each aggregate represents occupy adjacent positions in every consumer’s preferences, then RU-rationality implies ARU-rationality regardless of the composition distribution. The second is menu-independent composition of $\lambda$ (Proposition 4.2): if the composition distribution does not vary across menus, then RU-rationality implies ARU-rationality regardless of the preference structure. These results provide practical guidance for constructing aggregates–alternatives should be grouped if they are close substitutes and if their availability does not vary systematically across markets.
@@ 1.2 Simulation results
Our results imply that when either condition—non-overlapping preferences or menu independence—fails, the observed choice frequencies diverge from those predicted by an ARUM. Estimating an ARUM (such as logit model over aggregates) can therefore generate substantial bias. Assessing this bias empirically is crucial for understanding the practical relevance of our theory.
We examine this prediction through simulations. We assume that the true choice dataset defined on the underlying alternatives $\mathscr{X}$ is a logit model with fixed utility levels. We then fix values of a composition distribution $\lambda$ and an aggregation correspondence $X$ and construct the reduced dataset $\rho$ over aggregates. Finally, we re-estimate utilities from $\rho$ as if it followed a logit model over aggregates. When $\lambda$ is menu-independent or preferences are non-overlapping, $\rho$ is consistent with an ARUM and bias must be low, otherwise the bias should become large as $\rho$ cannot be represented by an ARUM.
As our measure of bias, we examined the difference in estimated utilities between two atomic aggregates, and we also computed the geometric distance from $\rho$ to the set of ARUMs. Both measures increase as $\lambda$ becomes more menu-dependent or preferences more overlapping. In some cases, the bias was substantial enough to reverse the ranking of alternatives—for example, even when $u(x)>u(y)$ in the true model, the estimates produced $\hat{u}(y)>\hat{u}(x)$. These findings align with our theoretical predictions and underscore the empirical importance of potential biases. Further details are provided in Section 5.
@@ 1.3 Related Literature
There is no paper in the stochastic-choice literature that directly studies aggregation when the composition of aggregates is unknown to the analyst and may vary across individuals and menus. There are, however, several related papers that share aspects of our motivation or mathematical setup.
@@
A first related line of research is the recent work on random attention (or limited consideration) models, such as Cattaneo et al., 2020; Aguiar et al., 2023; Brady and Rehbeck, 2016; Manzini and Mariotti, 2014; Horan, 2019. See Strzalecki, 2022 for a survey. These papers study environments in which agents do not evaluate the full menu but instead randomly consider a subset. In particular, Horan (2019) analyzes a setting in which the choice frequency of the outside option is not observed.
The resemblance, however, is limited. Random-attention models are driven by unobserved consideration sets, whereas our setting is driven by aggregation-induced unobservability: the analyst observes only aggregates, whose internal composition can differ across consumers and menus. This difference has sharp implications. In our framework, when an aggregate is present, at least one of its underlying components is available; random-attention models impose no such requirement. Moreover, in the random attention model, it is assumed that the analyst observes an agent’s choices over underlying alternatives; and they violate RU rationality because some options are not considered. In contrast, we maintain random utility at the underlying level; violations arise only at the aggregated level.
@@
Azrieli and Rehbeck (2022) studies a mathematically related model but with a different focus. They characterize marginal stochastic choice—frequencies not conditioned on particular menus—assuming the analyst observes the distribution of menus but not within-menu choices. By contrast, in our framework the distribution over available alternatives is unobserved and captured by $\lambda$, while we observe reduced stochastic choice over aggregates. Although Azrieli and Rehbeck (2022) also allows for endogenous menu availability, their analysis centers on models such as Gul and Pesendorfer (2001) and Kreps (1979), and addresses rationalizability rather than estimation bias. Apesteguia and Ballester (2016) also study aggregation of stochastic choice data, but with a different motivation. They examine whether aggregated choice behavior, obtained by combining individual choices, remains in the same class as the underlying individual behaviors.
@@
Kono et al. (2025) also studies the RUM with an outside option, but in a different setup. The paper focuses on the outside option case. Moreover, from the perspective of the observed dataset, the key distinction between the two papers is that the present study takes the reduced dataset $\rho$ as the observed object, whereas Kono et al. (2025) treats the full dataset over the underlying alternatives (with missing information) as given. In addition, Kono et al. (2025) assumes the outside option always corresponds to the full set of underlying alternatives—that is, $\lambda$ is menu-independent and degenerate. Even in this idealized case, much of the implications of the RUM are lost. In the present paper, we extend that framework by allowing $\lambda$ to be non-degenerate or menu-dependent, and we characterize the observable implications of the RUM under this more general specification.
@@
A few papers in the empirical IO literature such as Brownstone and Li (2017), Stafford (2018), and Huang and Rojas (2013) have examined potential biases arising from aggregation and misspecification of the outside option. More recently, Zhang (2023) analyzed biases resulting from misspecifying market size, and hence the market share of the outside good. However, all of these papers are purely empirical in nature and do not address the issue of unknown composition or outside options theoretically.
@@
Although less directly related, classical statistical work has developed a comprehensive framework for making inferences from data that are partially observed or categorized. For example, see Heitjan (1989) for a review of grouped data; Blumenthal (1968) and Nordheim (1984) for categorized data; and Heitjan and Rubin (1990, 1991) for coarse data. In this framework, subsets of the sample space are observed rather than exact realizations. [^4] For example, if the sample space is the real line (e.g., blood pressure), we may only observe a random variable taking interval values (e.g., an interval of blood pressures containing the true value). A key distinction from our setting is that, in our model, uncertainty arises from the actual underlying alternatives that an agent faces, rather than from the data generating process. We model decision making of consumers under the random utility model, which does not appear in this statistical literature.
@@ 2 Setup
Let ${\mathscr{A}}=\{a_{0},....,a_{N}\}$ be the finite set of aggregates. There are two types of aggregates; one is atomic aggregate, which has only one underlying alternative. The other one is a non-atomic aggregate such as the outside option, which has multiple different underlying alternatives. (See Example 2.1 below.) When there is no danger of confusion, we simply say aggregates rather than non-atomic aggregates. We denote the set of atomic aggregates by $\mathscr{A}_{A}$; denote the set of non-atomic aggregates by $\mathscr{A}_{N}$. Thus, $\{\mathscr{A}_{A},\mathscr{A}_{N}\}$ forms a partition of $\mathscr{A}$. [^5] That is, $\mathscr{A}_{A}\cup\mathscr{A}_{N}=\mathscr{A}$ and $\mathscr{A}_{A}\cap\mathscr{A}_{N}=\emptyset$.
Let $\mathscr{X}$ be the set of underlying alternatives. We assume that the underlying set $\mathscr{X}$ is finite. [^6] We will make an assumption on the richness of $\mathscr{X}$ later for some results. In this paper, the partition $\{\mathscr{A}_{A},\mathscr{A}_{N}\}$ and $\mathscr{X}$ are exogenously given.
We now define an aggregation correspondence, which connects $\mathscr{A}$ and $\mathscr{X}$.
@@
A function $X:\mathscr{A}\to 2^{\mathscr{X}}\setminus\{\emptyset\}$ is called an aggregation correspondence if (i) $X(a_{i})\cap X(a_{j})=\emptyset$ for any distinct $i,j$; (ii) $X(a)$ is singleton if $a\in\mathscr{A}_{A}$; (iii) $X(a)$ is non-singleton if $a\in\mathscr{A}_{N}$.
We call $X$ an aggregation correspondence because the correspondence (or more precisely, its inverse) specifies how each underlying alternative is assigned to an aggregate. Condition (i) requires that every underlying alternative be assigned to exactly one aggregate. Conditions (ii) and (iii) justify the interpretation that $\mathscr{A}_{A}$ represents the set of atomic aggregates, while $\mathscr{A}_{N}$ represents the set of non-atomic aggregates.
In many parts of the paper, we consider the setup in which there is only one non-atomic aggregate $a_{0}$ called outside option. In the setup, we assume $\mathscr{A}_{N}=\{a_{0}\}$.
@@
(Cereal Choice) Consider a similar setup to Nevo (2001), who estimates demand over the popular cereal brands. As mentioned in the introduction, a key feature of the analysis is the inclusion of an outside good, representing all breakfast options not explicitly listed among the popular cereal brands in the dataset. In the setup, all available alternatives are